substitution
x+2y=13
3x-5y=6

Answer by checkley71(8403)

X+2Y=13 OR X=13-2Y
3(13-2Y)-5Y=6
39-6Y-5Y=6
-11Y=6-39
-11Y=-33
Y=-33/-11
Y=3 ANSWER.
X=13-2*3
X=13-6
X=77 ANSWER.
PROOF
7+2*3=13
7+6=13
13=13